OpenStudy (anonymous):

Find the equation of the straight line that: passes through (1 , 4) and is parallel to y = 3x Anyone can help with this? Thanks

OpenStudy (anonymous):

Parallel lines have the same slopes, so you have y=3x + c So: 4 = 3(1) + c 4 = 3 + c 1 = c Therefore y=3x +1